WHAT IS THE ACCELERATION DUE TO GRAVITY AT A GRAVITY AT A DISTANCE 3r FROM THE CENTER OF THE EARTH IF THE GRAVITATIONAL POTENTIAL AT A  DISTANCE r FROM THE CENTER OF THE EARTH IS v ? [ assume  r>R , R= radius of earth ]

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Gravitational potential at a distancerfrom center of Earth=V=\dfrac{GM}{r}

Acceleration due to gravity at distance3rfrom center of earth=\dfrac{GM}{(3r)^2}=\dfrac{GM}{9r^2}=\dfrac{V}{9r}